OpenStudy (anonymous):

find the difference quotient { f(x+h)-f(x) } with a denominator of h. h is not equal to 0, for the function of f(x)=4x^2+6x

OpenStudy (anonymous):

\[\frac{f(x+h)-f(x)}{h}=\frac{4(x+h)^2+6(x+h)-(4x^2+6x)}{h}\]

OpenStudy (anonymous):

\[=\frac{4(x^2+2xh+h^2)+6x+6h-4x^2-6x}{h}\]

OpenStudy (anonymous):

\[=\frac{4x^2+8xh+4h^2+6x+6h-4x^2-6x}|{h}\]

OpenStudy (anonymous):

=\[\frac{8xh+4h^2+6h}{h}\]

OpenStudy (anonymous):

\[=\frac{h(8x+6+4h}{h}\]

OpenStudy (anonymous):

\[=8x+6+4h\]

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

wow thank you. ur a guinness.. yes helps a lot to see it broken down

OpenStudy (anonymous):

thnx